OpenStudy (anonymous):

How do i find the area of the rectangle 7x^2 6x^4 would I do this 7*6=42 42x add the exponents 2+4=6 42x^6????

OpenStudy (anonymous):

area of square formula A = l w Exponent rules also apply here

OpenStudy (anonymous):

??? im confused

OpenStudy (anonymous):

EXAMPLE NOT YOUR MATH PROBLEM AREA OF RECTANGLE 9 X^4 8 X^7 ( 9 X ^ 8 )( 8 X ^ 7) 9 * 8 = 72 EXPONENT RULE X ^ M+N= (X ^ M ) (X ^ N)

OpenStudy (skullpatrol):

Yes your answer is correct: would I do this 7*6=42 42(x^2)(x^6) add the exponents 2+4=6 42x^6????